Circular No. 9135

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET P/2009 WX_{51} (CATALINA) A. Mainzer, Jet Propulsion Laboratory, reports that WISE 12- and 24-

μm images obtained on Apr. 2 and 3 of 2009 WX_{51}, an apparently asteroidal object found by the Catalina Sky Survey (cf. MPEC 2009-W76; discovery observation tabulated below), show a 25′′ coma and a bright, fan-shaped tail extending ∼ 50′′ beyond the coma spanning from west-southwest to west-northwest — appearing as bright as P/2010 B2 was on the WISE discovery images of that comet (IAUC 9115; so perhaps magnitude V ∼ 20); unfortunately, there are no ground-based observations of 2009 WX_{51} since 2009 Dec. 20 (cf. MPS 305307, 306814, 307444, and 308762).

2009 UT α_{2000} δ_{2000} Mag. Observer
Nov. 22.24934 4^{h}25^{m}19.^{s}84 +40^{o}27′43′′.2 18.4 Catalina

The WISE astrometry, the following improved orbital elements, and an ephemeris appear on MPEC 2010-G105.

Epoch = 2010 Feb. 13.0 TT

T = 2010 Jan. 31.0741 TT  ω = 118.^{o}0312
e = 0.740286  Ω = 31.7417 }2000.0
q = 0.799956 AU  i = 9.5914

a = 3.080147 AU n^{o} = 0.1823253 P = 5.41 years

COMET C/2007 Q3 (SIDING SPRING) F. Colas, Observatoire de Paris, forwards a report by F. Manzini (Soz-

zago, Italy) of an 18th-magnitude condensed fragment in the tail of this comet, along with a fainter fragment, on Mar. 13.0 UT; follow-up images by Colas on Mar. 14.1 using the Pic du Midi 1-m telescope confirmed a single 18th-mag object some 6′′ from the main nuclear condensation in p.a. 240^{o}. N. J. Howes, Cherhill, Wiltshire, U.K., reported a fragment of mag ∼ 18 located ∼ 3′′ in p.a. 240^{o}–253^{o} from the primary component on Mar. 17 on images taken through three filters with the Faulkes Telescope North; his images from Apr. 12 fail to show component ‘B’, suggesting its disintegration. E. Bryssinck, Kruibeke, Belgium, found a secondary component 5′′.5 from the nuclear condensation in p.a. 258^{o} in eight images (resolution 0′′.5/pixel) taken on Mar. 19.397 remotely with a 0.4-m reflector at the Tzec Maun Observatory in New Mexico. Reports suggest that component ‘B’ was ∼ 4 mag fainter than primary component ‘A’. The astrometry of this new component ‘B’ from Mar. 17 to Apr. 9 has been included on MPECs 2010-F87, 2010-G14, and 2010-G70.

2010 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green